OpenStudy (anonymous):

Will medal and fan. What is the x-intercept of the following line? 4x -7y = -28

OpenStudy (anonymous):

The x-intercept is what 'x' is when y is 0. So we can plug in 0 for y. 4x - 7y = -28 4x - 7(0) = -28 4x - 0 = -28 4x = -28 Now divide 4 to both sides, what's -28 / 4?

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

Yes, so the x-intercept is -7.

OpenStudy (anonymous):

(: